OpenStudy (anonymous):

What is the volume of the solid generated by revolving the region bounded by the given lines and curves about the x-axis? y=sq root(2x+3) y=0, x=0, x=1

OpenStudy (anonymous):

..

OpenStudy (dumbcow):

 \[V = \pi \int\limits_{0}^{1}(\sqrt{2x+3})^{2} dx\]